Suppose
m∠A = (2x + 4)°
and
m∠B = (3x − 34)°.
Find the measure of each angle (in degrees) of
▱ABCD.
m∠A =
°
m∠B =
°
m∠C =
°
m∠D =
°

Answer by ikleyn(52787)

.

In each parallelogram, the adjacent angles are supplementary, i.e. sum up to 180 degrees.


Angles A and B are adjacent, so you write this equation

    (2x + 4 ) + (3x - 34) = 180  degrees.


Simplify and find x

     2x + 4 + 3x - 34 = 180

         5x  - 30     = 180

             5x       = 180 + 30

             5x       = 210

              x       = 210/5 = 42.


Having x, find the angles.  Use the given formulas.

Solved (which means that the way is shown and you are fully instructed).


-------------------


Comment from student : ok so since i have x how would i find the measures of a,b,c and d ? should i divide 42 into 180?


My response : In your post, there are formulas expressing angles via x.

Substitute the value of x= 42 degrees in these formulas and calculate the angles.

Keep in mind that in parallelogram opposite angles are congruent, i.e. have equal degree measures.

It will help you to find all angles.